OpenStudy (anonymous):

/Users/westonpinyan/Desktop/Screen Shot 2013-10-31 at 7.15.48 PM.png

OpenStudy (anonymous):

Will someone help me on this one question on this URL?

OpenStudy (jdoe0001):

as soon as we can sit by your side and use your keyboard, yes

OpenStudy (jdoe0001):

attach the file here, using the [Attach File] blue button

OpenStudy (jdoe0001):

test which 2 shorter sides, ADD UP to the longest side in the triangle

OpenStudy (anonymous):

I know that, I have worked out both of them and nether of them is a perfect square.

OpenStudy (jdoe0001):

? what do you mean?

OpenStudy (anonymous):

a2+b2=c2. I have used that equation and a and b does not match up with c

OpenStudy (jdoe0001):

hmm the non-squared version is not supposed to match, you're correct, the squared version is 9 units means, the side is 3 long 36 units means the side is 6 long 64 units means the side is 8 long \(\bf 3^2+6^2=8^2\implies 9+36\ne 64\)

OpenStudy (jdoe0001):

let's check the 2nd one 25 units, 5 long 144 units, means is 12 long 169 units, means is 13 long \(\bf 5^2+12^2=13^2\implies 25+144=169\)